OpenStudy (anonymous):

The larger of two numbers is 6 times the smaller. Their sum is 147. Find the numbers.

OpenStudy (anonymous):

okay if we call one number x and the other y, what can you tell me about the relationship between x and y?

OpenStudy (anonymous):

they have to come to 147

OpenStudy (anonymous):

okay, the sum of x and y is 147. so can you write that sentence in mathematical form?

OpenStudy (anonymous):

x+y=147 y x 6=x

OpenStudy (anonymous):

very good. so, x = 6y but we know that x+y =147, so 6y+y=147

OpenStudy (anonymous):

x=67.5 y-79.5?

OpenStudy (anonymous):

no, how did you get that?

OpenStudy (anonymous):

idk am so lost

OpenStudy (anonymous):

6y+y = 147. how do you get y from that equation?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Im lost why Im here trying to figure it out

OpenStudy (anonymous):

6y+y = 7y. so 7y =147. divide both sides by 7. y = 147/7 y =21

OpenStudy (anonymous):

but it is asking for 2 numbers would 21 be the smaller # or larger?